OpenStudy (anonymous):

Find the area of a segment formed by a chord 8" long in a circle with radius of 8".

OpenStudy (anonymous):

isnt the chord the same thing as the radius?

OpenStudy (anonymous):

You have to solve for the area of the sector and the area of the triangle area of the sector = pi * r^2 * (angle / 360) You need to determine the angle. You can calculate this via trigonometry. However, take note that the triangle bound by the chord and the radii is an equilateral triangle. Hence, the angle in this problem is 60 degrees. area of the triangle = 1/2 bh * b = the length of the chord; h = the height of the triangle, calculate this using trigonometry. You now have everything you need to solve the area of the segment.

OpenStudy (anonymous):

thank you :D